# Towards foundational LiDAR world models with efficient latent flow matching

### v1

@@ A.1 Training Setup
For all flow matching based generative (foundational) model training, different from previous methods gu2024dome ; li2024uniscene that used several thousands of training epochs, we used 4x RTX 4090 to train models for 200 epochs with batch size 8. We trained the VAE part for 100 epochs with batch size 16 if not specified.
Following prior workzhang2023copilot4d , we also adopt the AdamW as the optimizer with $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.99 for Flow matching training, 0.99 and 0.999 for VAE training. We set the weight decay of all normalization layers to 0 and all of other layers’ to 0.001. The learning rate schedule has a linear warmup followed by cosine decay (with the minimum of the cosine decay set to be 20% of the peak learning rate). We also use EMA with 0.9999 decay rate to ensure the updates of parameters are stable. During the training, we found that the conditional free guidance is also important in a flow-matching based framework. We randomly set 25% of historical latent in a batch to 0 during training to let model learn to generate future latents without any condition. In the sampling phase, we use a typical fuse method as shown in Eq. 3 to get the final output when it’s the $t$-th step, where s set to 2.
@@
Finally, we also noticed that the matching of noise scale and latent scale is important. Different from SD3’s VAE, the value of latent obtained from our VAE is actually smaller, with a standard deviation of about 0.02, which makes it easy to drown the signal in noise if we use standard Gaussian noise if we use it in the training phase. Therefore, we scaled the compressed latent by a factor of 10 in the implementation.
A.2 Latent space similarity metrics
@@
In Section 4.3, we introduce CKA(Centered Kernel Alignment) kornblith2019similarityneuralnetworkrepresentations and CKNNA(Centered Kernel Nearest-Neighbor Alignment )huh2024platonic to evaluate the similarity of latent space, both of them are insensitive to linear transformations. Specifically, we treat each latent pixel (location) as an individual sample i.e., every sample in CKA/CKNNA evaluation can be represented as $\bm{z}_{i}\in\mathbb{R}^{1\times c}$. For non-semantic to semantic adaptation subtask, given $\mathbf{k}$ samples from $\mathbf{e}$ latent, the representation metrics from $\bm{o}_{d}$ and $\bm{o}_{s}$ can be denoted by $\bm{X}\in\mathbb{R}^{u\times c}$ and $\bm{Y}\in\mathbb{R}^{u\times c}$ respectively. We first construct kernel matrix:
@@
where $\bm{K}_{c}$ and $\bm{L}_{c}$ are centered $\bm{K}$ and $\bm{L}$, using $\bm{K}_{c}=HKH$, $H=I_{n}-\frac{1}{n}\bm{1}_{n}\bm{1}_{n}^{T}$. $\bm{1}_{n}$ is all-ones column vector. CKNNA is a revised version of CKA which focus on the local manifold similarity.
@@
A.3 Inception-based metrics
@@
To obtain latent features from generated and ground-truth samples under comparable conditions, we retrain an autoencoder on $\bm{o}_{s}$. The autoencoder is based on the MinkowskiUNet32 architecture choy20194d , but we adapt it to our 200 × 200 × 16 occupancy inputs by reducing the feature channels from {64, 128, 256, 512} with four down-sampling stages to {32, 64, 128} with three down-sampling stages.
For the output of 6 frames of future semantic occupancy, we evaluated frame-wise FID and KID. Specifically, we first extracted all non-empty voxels from the downsampled semantic occupancy (size 25, 25, 2) and took the average, rather than performing global pooling as in previous work, to avoid the influence of empty voxels on the metrics. Given the flattened feature $\bm{z}^{g}_{i}$ and $\bm{z}^{e}_{j}$ from single sample, we use $\bm{\mu}^{g}$, $\bm{\mu}^{e}$ and $\bm{\Sigma}^{g}$, $\bm{\Sigma}^{e}$ to represent the mean and channel co-variance matrix over all samples M ground truth samples and N estimated samples. Then we calculate FID and with the following equations.
@@
For KID, we use a RBF kernel based unbiased-statistic estimation version as shown in equation 10.
@@
where $\phantom{=}\text{where }k_{\sigma}(\bm{x},\bm{y})=\exp\!\Bigl{(}-\lVert\bm{x}-\bm{y}\rVert^{2}/(2\sigma^{2})\Bigr{)}.$
A.4 Negative log likelihood evaluation
A.4.1 Computation of exact log probability density
Due to the stochastic nature of our proposed conditional flow matching model, pairwise-sample-comparison metrics such as IoU and mIoU provide a limited measure of model quality, since these “deterministic" metrics penalize models from generating diverse but plausible predictions of the future that are substantially different from the recorded future. In other words, the model’s ability to capture the uncertainty in future prediction is not measured well by IOU and mIOU. Therefore, we also evaluate the exact log probability of our CFM model that doesn’t penalize it substantially for assigning probability density to other modes.
With our CFM model, $G_{\theta}$, we can compute the log probability of any generated future prediction or future ground-truth samples with the following methods. First, as described by previous works chen2019neuralordinarydifferentialequations , the generative process of a continuous normalizing flows works as the following: starting from a sample from a base distribution $\bm{z}_{0}\in p_{z_{0}}(z_{0})$ and a parametrized ODE $G_{\theta}=G(z(t),t;\theta)$ which is the flow function, we can obtain $\bm{z}(t_{1})$ from the target distribution by solving for the initial value problem $\bm{z}(t_{0})=z_{0}$, $\frac{\partial z(t)}{\partial t}=G$. In this process, the rate of change of log-probability-density follows the instantaneous change of variables formula chen2019neuralordinarydifferentialequations :
@@ A.4.1 Computation of exact log probability density
[…]
@@
where D is the number of scalar dimensions of the sample. This enables us to determine exact log likelihood of any generated sample or sample from the future occupancy latent space (any sample from future occupancy distribution) numerically.
A.4.2 Details on NLL evaluation
For all log likelihood values computed, we use an Euler ODE solver with a step size of 0.02, relative and absolute tolerance of $1\times 10^{-5}$. Trace of Jacobian is obtained by utilizing Hutchinson’s trace estimator with the probe vector sampled from a Rademacher distribution.
@@ A.4.2 Details on NLL evaluation
where D is the number of scalar dimensions of the latent sample x and the division by $\ln 2$ converts the unit from nats to bits.
A.4.3 DDPM log likelihood computation
For discrete DDPM-based model, we need to construct and solve the IVP introduced previously with mathematically-equivalent representations for the flow function $G$.
@@ A.4.3 DDPM log likelihood computation
A.4.4 Results on semantic occupancy forecasting task
The NLL values in terms of BPD and the corresponding standard deviation (across entire validation set; we use ground-truth samples not the generated predictions) obtained on semantic occupancy forecasting task is shown in table 5. Our semantic occupancy forecasting model conditioned on historical/past trajectory achieves comparable NLL values with DOME gu2024dome : only 0.25% absolute difference (4.1% relative) while DOME utilizes future trajectory in the condition. For our model conditioned on future trajectory, the NLL value is lower than that of DOME by 25% relative difference (1.49% absolute). This shows that our CFM model better fits the true future-occupancy distribution than the previous (stochastic) SOTA model.
In terms of standard deviation in NLL, for our models (conditioned on either history or future trajectory), the standard deviation is at least 86% (relative) less than that of DOME. This statistical metric further support the conclusion drawn above by demonstrating the high consistency of our model’s performance in terms of assigning high probability to correct futures.
Table 5: Mean negative log-likelihood for different semantic occupancy forecasting models.
@@ A.4.4 Results on semantic occupancy forecasting task
Different from image, LiDAR occupancy from a BEV perspective has a clear depth correspondence between grid cells and environments. However, the global pooling operation will take the average of features in all grid cells, agnostic to geometric location. Here, inspired by ran2024towards , we also measured the FID and MMD of feature concatenated by features from different depth bins. Specifically, in the obtained $\bm{z}_{i}^{e}\in\mathbb{R}^{25\times 25\times 2}$, every cell stand for $3.2m\times 3.2m$ area in the environments. Based on the distance from the center point, we divide the area into three zones: within ±8 m, ±8 to ±24 m, and ±24 to ±40 m. We then take the average of the non-empty voxels in each zone and concatenate them to get the final feature for evaluation, denoted as $\text{FID}_{r}$ and $\text{KID}_{r}$ in Table 7. The original FID and KID results are shown in Table 6.
Table 6: FID and KID(×$\times 10^{-2}$) comparison for our method and other generative-based and deterministic-based models.
Method 1s 2s 3s AVG
FID KID FID KID FID KID FID KID
OccWorld 9.54 11.7 8.56 10.46 7.78 9.80 8.62 10.60
DOME 4.37 4.39 5.09 5.44 5.63 6.53 5.03 5.42
Ours(Hist. traj) 1.67 1.49 2.81 2.85 4.02 4.74 2.83 3.02
Ours(Fut. traj) 1.61 1.47 2.90 3.05 3.91 4.55 2.80 3.02
Table 7: $\text{FID}_{r}$ and $\text{KID}_{r}$(×$\times 10^{-2}$) comparison for our method and other generative-based and deterministic-based models.
@@
Interestingly, we observed that the FID/KID of OccWorld slightly decrease (around 2%) when we forecast longer-term future results. We contribute this to the auto-regressive displacement forecasting design in OccWorld, which preserve more diversity when forecasting longer term, although the overall performance is lower. Specifically, we calculated the category variance at different time horizons(all numbers follow in ×$\times 10^{-3}$): for the ground truth, the variance fluctuated between 3.055 and 3.021, while in occworld it rose from 2.865 at t=0.5 seconds to 3.038, and our model fell from 3.086 to 2.821.
In Table 8, we report the future occupancy forecasting performance and computational efficiency comparison (include parameter count). It is shown that our proposed framework achieves state-of-the-art performance while being more efficient across all efficiency metrics. Our model outperforms OccWorld and DynamicCity while having half the computational cost, less than 66% of parameters, and 1.1 times higher in FPS. Compared to DOME, with future trajectory information, our model achieves at least 5.5% better absolute mIoU while only using 4.38% of flops and 15 times less parameters.
[…]
OccWorld zheng2024occworld 17.13 26.63 1347.09 72.39 2.56
RenderWorld yan2024renderworld 20.80 30.08 416
DynamicCity bian2024dynamiccity 21.56 29.97 774.44 45.43 19.30
Ours 23.33 31.78 389.46 30.37 22.22
DOME gu2024dome 22.18 32.13 8891.98 444.07 0.76
Ours 28.49 37.52 389.46 30.37 21.43
@@ Appendix B Dataset introduction
As mentioned earlier, we used nuScenes Semantic Occupancy($\bm{o}_{s}$), KITTI360 raw sweepsliao2022kitti , and our own collection of indoor LiDAR data to test the three subtasks. In semantic occupancy forecasting task, we further divide nuScenes Semantic Occupancy to pretrain set ($\bm{o}_{s}^{\prime}$) and fine-tuning set. For pretrain, we use voxelized lidar sweeps in nuScenes dataset. In Table 9, we list the specific number of frames of all these datasets after down-sampling to 2Hz.
@@
Specificially, for indoor LiDAR dataset, it consists of 25 sequences that we collected in 2 different scenarios. The indoor data was collected using an OSO-128 LiDAR sensor with a 90 degree vertical field of view. A robot, modelled with unicycle kinematics, was manually driven through a university building over the course of five days. An expert human operator controlled the robot, skillfully avoiding pedestrians while navigating toward designated goals. As shown in Figure 5, compared to the data in autonomous driving scenario, where the vast majority of vehicles move parallel to ego vehicles, the data we collected include more complex trajectories with more objects (people).
@@
As we mentioned in the main text, the amount of data after downsampling is still large considering that we only collected this data in 2 scenes. This makes it easier for the model to learn domain-specific knowledge, and explains why the performance of our pre-trained model drops after using more than 50% of the data.
@@ Appendix C Detailed pretraining and fine-tuning results
In this section, we explain more details about the results from our experiments on how different pretrained/from-scratch VAE and CFM combinations perform on each of the three subtasks, which, again, are high-beam LiDAR adaptation, indoor occupancy forecasting, and semantic occupancy forecasting. For LoRAhu2021loralowrankadaptationlarge , across all our downstream tasks, we choose rank of 128 and $\alpha$ of 128. Please check section E for details on this selection. All models present in this section are trained for 40 epochs.
@@ C.1 High-beam adaptation and indoor occupancy forecasting
In Table 10 and Table 11, we organize the performance of each of the pretrain/from-scratch combinations (VAE pretrained and CFM pretrained and fine-tuned with LoRAhu2021loralowrankadaptationlarge , both pretrained and full-parameter fine-tuned, only CFM pretrained, and both from-scratch) for high-beam adaptation and indoor occupancy forecasting in terms of IoU. IoU is computed through inferring the trained CFM model with NFE=10. Justification on this decision of this value of NFE is shown in section F.4. For high-beam adaptation task, pretrained CFM model is trained on original LiDAR sweeps, denoted by $G_{\theta}^{v}$. From the result, it is evident that full-parameter pretrained (both or only-CFM, with or without LoRA) yields exceeding performance over from-scratch trainings. Specifically, At 10% of fine-tuning data usage, our pretrained model with full-parameter fine-tuning achieves more than 14.48% relative performance improvement (more than 1.5% absolute performance improvement) across the three-second forecasting horizon and even at 100% data usage, our model still offers an absolute performance growth of about 1%. This demonstrates the effectiveness of our pretrained foundational in transferring geometric knowledge learned from 32 beams LiDAR data to tasks that demands different geometric properties (64 beams in this case) at different availability of fine-tuning data.
For indoor occupancy forecasting task, we use the same pretrained CFM model as the high-beam adaptation task. From the result, it is shown that at low fine-tuning data availability (10% and 25%), our pretrained models achieve exceeding performance over the from-scratch training. But when the amount of fine-tuning data increases to more than 50%, the from-scratch training obtains comparable or even better performance. Again, we explain this performance threshold to be resulting from the fine-tuning data amount over-rides the pretraining data and the lack of variance in the geometry of indoor occupancy dataset collected.

### v2

@@ A.1 Training Setup
For all flow matching based generative (foundational) model training, different from previous methods gu2024dome ; li2024uniscene that used several thousand training epochs, we used 4x RTX 4090 to train models for 200 epochs with a batch size of 8. We trained the VAE part for 100 epochs with a batch size of 16 if not otherwise specified.
Following prior workzhang2023copilot4d , we also adopt the AdamW as the optimizer with $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.99 for flow matching training, 0.99 and 0.999 for VAE training. We set the weight decay of all normalization layers to 0 and all of the other layers’ to 0.001. The learning rate schedule has a linear warmup followed by cosine decay (with the minimum of the cosine decay set to be 20% of the peak learning rate). We also use EMA with a 0.9999 decay rate to ensure the updates of parameters are stable. During the training, we found that the classifier free guidance is also important in a flow-matching based framework. We randomly set 25% of historical latents in a batch to 0 during training to let the model learn to generate future latents without any conditions. In the sampling phase, we use a typical fuse method as shown in Eq. 3 to get the final output when it’s the $t$-th step, where s is set to 2.
@@
Finally, we also noticed that the matching of noise scale and latent scale is important. Different from SD3’s VAE, the value of latent obtained from our VAE is actually smaller, with a standard deviation of about 0.02, which makes it easy to drown the signal in noise if we use standard Gaussian noise in the training phase. Therefore, we scaled the compressed latent up by a factor of 10 in the implementation.
A.2 Evaluation metrics
In this section, we will provide the details of all metrics used to further evaluate the performance of our proposed pipeline. All results will be presented in the last part of this section.
A.2.1 Latent space similarity metrics
@@
In Section 4.3, we introduce CKA (Centered Kernel Alignment) kornblith2019similarityneuralnetworkrepresentations and CKNNA(Centered Kernel Nearest-Neighbor Alignment )huh2024platonic to evaluate the similarity of latent spaces as both metrics are insensitive to linear transformations. Specifically, we treat each latent pixel (location) as an individual sample (i.e., every sample in CKA/CKNNA evaluation can be represented as $\bm{z}_{i}\in\mathbb{R}^{1\times c}$). For the non-semantic to semantic adaptation subtask, given $\mathbf{k}$ samples from $\mathbf{e}$ latent, the representation metrics from $\bm{o}_{d}$ and $\bm{o}_{s}$ can be denoted by $\bm{X}\in\mathbb{R}^{u\times c}$ and $\bm{Y}\in\mathbb{R}^{u\times c}$ respectively. We first construct the kernel matrix:
@@
where $\bm{K}_{c}$ and $\bm{L}_{c}$ are centered $\bm{K}$ and $\bm{L}$, using $\bm{K}_{c}=HKH$, $H=I_{n}-\frac{1}{n}\bm{1}_{n}\bm{1}_{n}^{T}$. $\bm{1}_{n}$ is an all-ones column vector. CKNNA is a revised version of CKA that focuses on the local manifold similarity.
@@
A.2.2 Inception-based metrics
Fidelity measurement
@@ A.2.2 Inception-based metrics
To obtain latent features from generated and ground-truth samples under comparable conditions, we retrain an autoencoder on $\bm{o}_{s}$. The autoencoder is based on the MinkowskiUNet32 architecture (sparse)choy20194d , but we adapt it to our 200 × 200 × 16 occupancy inputs by reducing the feature channels from {64, 128, 256, 512} with four down-sampling stages to {32, 64, 128} with three down-sampling stages. The training objective is to recover the compressed latent as much as possible under a cross-entropy loss.
For the output of 6 frames of future semantic occupancy, we evaluated frame-wise FID and KID. Specifically, we first extracted all non-empty voxels from the down-sampled semantic occupancy (size 25, 25, 2) and took the average, rather than performing global pooling as in previous work, to avoid the influence of empty voxels on the metrics. Given the flattened feature $\bm{z}^{g}_{i}$ and $\bm{z}^{e}_{j}$ from individual samples, we use $\bm{\mu}^{g}$, $\bm{\mu}^{e}$ and $\bm{\Sigma}^{g}$, $\bm{\Sigma}^{e}$ to represent the mean and channel covariance matrix over M ground truth samples and N estimated samples. Then we calculate FID with the following equations.
@@
For KID, we use a RBF-kernel-based unbiased-statistic estimation version as shown in equation 10.
@@
$\phantom{=}\text{where }k_{\sigma}(\bm{x},\bm{y})=\exp\!\Bigl(-\lVert\bm{x}-\bm{y}\rVert^{2}/(2\sigma^{2})\Bigr).$
Temporal Consistency Assessment
FID and KID provide assessments of single-frame fidelity, whilst for multi-frame continuity we adhere to previous methodology by employing FVD for evaluation. Given that no existing RGB video encoder can capture this 4D temporal information, we modified the proposed VAE to enable it to measure the continuity of 4D occupancy. Within the encoder and decoder shown in Figure 2, we respectively incorporated 3D attention/conv layers to capture temporal information.
A.2.3 Negative log likelihood evaluation
Computation of exact log probability density
Due to the stochastic nature of our proposed conditional flow matching model, pairwise-sample-comparison metrics such as IoU and mIoU provide a limited measure of model quality, since these "deterministic" metrics penalize models for generating diverse but plausible predictions of the future that are substantially different from the recorded future. In other words, the model’s ability to capture the uncertainty in future prediction is not measured well by the IoU and mIoU metrics. Therefore, we also evaluate the exact log probability of our CFM model that doesn’t penalize it substantially for assigning probability density to other modes.
With our CFM model, $G_{\theta}$, we can compute the log probability of any generated future prediction or future ground-truth samples with the following methods. First, as described by previous works chen2019neuralordinarydifferentialequations , the generative process of a continuous normalizing flow works as described below: starting from a sample from a base distribution $\bm{z}_{0}\in p_{z_{0}}(z_{0})$ and a parametrized ODE $G_{\theta}=G(z(t),t;\theta)$ which is the flow function, we can obtain $\bm{z}(t_{1})$ from the target distribution by solving for the initial value problem $\bm{z}(t_{0})=z_{0}$, $\frac{\partial z(t)}{\partial t}=G$. In this process, the rate of change of log-probability density follows the instantaneous change of variables formula chen2019neuralordinarydifferentialequations :
@@ A.2.3 Negative log likelihood evaluation
$\begin{cases}z_{0}=\int_{t_{1}}^{t_{0}}G(z(t),t;\theta)\,dt\\ \log p(x)-\log p_{z_{0}}(z_{0})=\int_{t_{1}}^{t_{0}}-Tr(\frac{\partial G}{\partial z(t)})\,dt\end{cases}$ (13)
@@
where D is the number of scalar dimensions of the sample. This enables us to determine the exact log likelihood of any generated sample or sample from the future occupancy latent space (any sample from the future occupancy distribution) numerically.
Details on NLL evaluation
For all log likelihood values computed, we use an Euler ODE solver with a step size of 0.02, relative and absolute tolerance of $1\times 10^{-5}$. The trace of the Jacobian is obtained by utilizing Hutchinson’s trace estimator with the probe vector sampled from a Rademacher distribution.
@@
where D is the number of scalar dimensions of the latent sample x and the division by $\ln 2$ converts the unit to bits.
DDPM log likelihood computation
For discrete DDPM-based model, we need to construct and solve the IVP introduced previously with a mathematically equivalent representation for the flow function $G$.
@@
A.3 More results on semantic occupancy forecasting task
The NLL values in terms of BPD and the corresponding standard deviation (across the entire validation set; we use ground-truth samples not the generated predictions) obtained from the semantic occupancy forecasting task are shown in table 8. Our semantic occupancy forecasting model conditioned on historical/past trajectory achieves comparable NLL values with DOME gu2024dome : only 0.25% absolute difference (4.1% relative) while DOME utilizes future trajectory in the condition. For our model conditioned on future trajectory, the NLL value is lower than that of DOME by 25% relative difference (1.49% absolute). This shows that our CFM model better fits the true future-occupancy distribution than the previous (stochastic) SOTA model.
In terms of standard deviation in NLL, for our models (conditioned on either history or future trajectory), the standard deviation is at least 86% (relative) less than that of DOME. This statistical metric further supports the conclusion drawn above by demonstrating the high consistency of our model’s performance in terms of assigning high probability to correct futures.
Table 8: Mean negative log‐likelihood for different semantic occupancy forecasting models.
@@ A.3 More results on semantic occupancy forecasting task
Different from image, LiDAR occupancy from a BEV perspective has a clear depth correspondence between grid cells and environments. However, the global pooling operation will take the average of features in all grid cells, agnostic to geometric location. Here, inspired by ran2024towards , we also measured the FID and MMD of features from different depth bins. Specifically, in the obtained $\bm{z}_{i}^{e}\in\mathbb{R}^{25\times 25\times 2}$, every cell stands for $3.2m\times 3.2m$ area in the environments. Based on the distance from the center point, we divide the area into three zones: within ±8 m, ±8 to ±24 m, and ±24 to ±40 m. We then take the average of the non-empty voxels in each zone and concatenate them to get the final feature for evaluation, denoted as $\text{FID}_{r}$ and $\text{KID}_{r}$ in Table 9.
Table 9: $\text{FID}_{r}$ and $\text{KID}_{r}$($\times 10^{-2}$) comparison for our method and other generative-based and deterministic-based models.
@@
Interestingly, we observed that the $FID_{r}/KID_{r}$ of OccWorld decreases slightly (around 2%) when we forecast longer-term future results. We attribute this to the auto-regressive displacement forecasting design in OccWorld, which preserves more diversity at longer horizon, although the overall performance is lower. Specifically, we calculated the category variance at different time horizons (all numbers follow in $\times 10^{-3}$): for the ground truth, the variance fluctuated between 3.055 and 3.021, while in Occworld, it rose from 2.865 at t=0.5 seconds to 3.038, and our model fell from 3.086 to 2.821.
@@ Appendix B Dataset introduction
As mentioned earlier, we used nuScenes Semantic Occupancy ($\bm{o}_{s}$), KITTI360 raw sweepsliao2022kitti , and our own collection of indoor LiDAR data to test the three subtasks. In the semantic occupancy forecasting task, we further divide nuScenes semantic occupancy data into pretrain set ($\bm{o}_{s}^{\prime}$) and fine-tuning set. For pretraining, we use voxelized LiDAR sweeps in the nuScenes dataset. In Table 10, we list the specific number of frames of all these datasets after down-sampling to 2Hz.
@@
Specifically, for the indoor LiDAR dataset, it consists of 25 sequences that we collected in 2 different scenarios, using an OSO-128 LiDAR sensor with a 90-degrees vertical field of view. A robot, modeled with unicycle kinematics, was manually driven through a university building over the course of five days. An expert human operator controlled the robot, skillfully avoiding pedestrians while navigating toward designated goals. As shown in Figure 6, compared to the data in the autonomous driving scenario, where the vast majority of vehicles move parallel to ego vehicles, the data we collected includes more complex trajectories with more objects (people).
@@
As we mentioned in the main text, the amount of data after downsampling is still large considering that we only collected this data in 2 scenes. This makes it easier for the model to learn domain-specific knowledge and explains why the performance of our pretrained model drops after using more than 50% of the data.
@@ Appendix C Detailed pretraining and fine-tuning results
In this section, we explain more details about the results from our experiments on how different pretrained/from-scratch VAE and CFM combinations perform on each of the three subtasks, which, again, are high-beam LiDAR adaptation, indoor occupancy forecasting, and semantic occupancy forecasting. For LoRAhu2021loralowrankadaptationlarge , across all our downstream tasks, we choose a rank of 128 and $\alpha$ of 128. Please check section E for details on this selection. All models present in this section are trained for 40 epochs.
@@ C.1 High-beam adaptation and indoor occupancy forecasting
In Table 11 and Table 12, we organize the performance of each of the pretrain/from-scratch combinations (VAE pretrained and CFM pretrained and fine-tuned with LoRAhu2021loralowrankadaptationlarge , both pretrained and full-parameter fine-tuned, only CFM pretrained, and both from-scratch) for high-beam adaptation and indoor occupancy forecasting in terms of IoU. IoU is computed through inferring the trained CFM model with NFE=10. For the high-beam adaptation task, the pretrained CFM model is trained on original LiDAR sweeps, denoted by $G_{\theta}^{v}$. From the result, it is evident that full-parameter pretrained (both or only-CFM, with or without LoRA) yields exceeding performance over from-scratch trainings. Specifically, at 10% of fine-tuning data usage, our pretrained model with full parameter fine-tuning achieves more than 14.48% relative performance improvement (more than 1.5% absolute performance improvement) across the three-second forecasting horizon and even at 100% data usage, our model still offers an absolute performance growth of about 1%. This demonstrates the effectiveness of our pretrained foundational models in transferring geometric knowledge learned from 32 beam LiDAR data to tasks that demand different geometric properties (64 beams in this case) at different availability of fine-tuning data.
For the indoor occupancy forecasting task, we use the same pretrained CFM model as the high-beam adaptation task. From the result, it is shown that at low data availability for fine-tuning (10% and 25%), our pretrained models achieve exceeding performance over the from-scratch training. But when the amount of fine-tuning data increases to more than 50%, the from-scratch training obtains comparable or even better performance. Again, we explain this performance threshold to be resulting from the fine-tuning data amount which overrides the pretraining data and the lack of variance in the geometry of the indoor occupancy dataset collected.